# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 5.2.1 Slice-based Annotation
In 3D modalities like CT and MRI, adjacent 2D slices often exhibit significant semantic redundancy. Consequently, annotating only the key slices of each sample can reduce annotation costs. Representativeness-based methods have been widely applied in this line of work. For instance, Zheng et al. (2020) utilized autoencoders to learn the semantic features of each slice, then selected and annotated key slices from axial, sagittal, and coronal planes with a strategy similar to RA. Specifically, they initially trained three 2D segmentation networks and one 3D segmentation network, where the inputs for the 2D networks are slices from different planes. These segmentation networks were used to generate four sets of pseudo-labels and subsequently to train the final 3D segmentation network. Results showed that this slice-based strategy outperforms uniform sampling. Building upon this method, Wu et al. (2022c) incorporated a self-attention module into the autoencoder to enhance slice-level feature learning. In recent years, uncertainty methods have been introduced for selecting key slices. In interactive segmentation of 3D medical images, Zhou et al. (2021a) and their subsequent work (Zhou et al., 2022) introduced a quality assessment module to provide a predicted average IoU score for each slice. They chose the slice with the lowest IoU score in each volume for the next round of interactive segmentation. In muscle segmentation of CT images, Hiasa et al. (2020) selected key slices and key regions. This work adopted clustering to select key slices and further selected regions with high uncertainty within each key slice for annotation.
@@ 5.2.2 One-shot Annotation
Currently, most AL works require multiple rounds of annotation. However, this setting may not be practical in medical image segmentation. Multi-round annotation requires physicians to be readily available for each round of labeling, which is unrealistic in practice. If physicians cannot complete the annotations on time, the AL process must be suspended. In contrast, one-shot annotation eliminates the need for multiple interactions with physicians. It also allows for selecting valuable samples in a single round, thus reducing time costs. Both one-shot annotation and cold-start AL aim to select the most optimal initial annotations. However, the former allows for a higher annotation budget and strictly limits the number of interactions with experts to just one. Most relevant works combine self-supervised features and specific sampling strategies to achieve one-shot annotation. For example, RA (Zheng et al., 2019) is one of the earliest works in one-shot AL for medical image segmentation. They applied the VAE feature and a representativeness strategy to select informative samples for annotation in one shot. RA performed excellently in segmenting gland of pathological images, whole-heart of MRI images, and fungal of electron microscopic images. Jin et al. (2022b) combined features of contrastive learning with farthest-first sampling to achieve one-shot annotation. The proposed method demonstrated effectiveness on the ISIC 2018 and lung segmentation datasets. Additionally, Jin et al. (2023b) utilized auto-encoding transformations for self-supervised feature learning. They selected and annotated samples with high density based on reachable distance.
@@ 5.2.3 Annotation Cost
Current AL works often assume equal annotation costs for each sample. Yet, this is not the case in medical image segmentation, where the time to annotate different samples can differ greatly. AL techniques can better support physicians by considering annotation costs (e.g., annotation time). Nevertheless, there is still limited research in this specific domain. In detecting intracranial hemorrhage of CT scans, Kuo et al. (2018) combined predictive disagreement with annotation time to select samples for annotation. Specifically, they adopted the Jensen-Shannon divergence to measure the disagreement between the outputs of multiple models. Annotation time for each sample was estimated by the length of the segmentation boundary and the number of connected components. In this work, AL was framed as a 0-1 knapsack problem, and dynamic programming is used to solve this problem for selecting informative samples. In brain structure segmentation, Atzeni et al. (2022) further considered the spatial relationships between multiple regions of interest to more accurately estimate the annotation cost. Moreover, the average Dice coefficient of previous rounds was used to predict the average Dice for current segmentation results. They selected and annotated regions that can maximize the average Dice.
5.3 Active Learning in Medical Image Reconstruction
AL can also be applied in medical image reconstruction. AL methods can help minimize the observations needed for modalities that require a long imaging time. This accelerates the imaging process and shortens the waiting period for patients. In this section, we’ll explore the application of AL in the reconstruction of MRI, CT, and electron microscopy.
@@ 5.3 Active Learning in Medical Image Reconstruction
In addition to MRI imaging, AL has been employed in CT reconstruction as illustrated by Wang et al. (2022a). They adaptively chose the scanning angles tailored to individual patients, leading to a reduction in both radiation exposure and scanning duration. In electron microscopy, Mi et al. (2020) initially enhanced low-resolution images to high-resolution and then predicted the location of region-of-interest and reconstruction error. A weighted DPP based on reconstruction error was applied to select pixels that needed rescanned. Results showed that weighted DPP maintained both low reconstruction error and spatial diversity.

### v3

@@ 5.2.1 Slice-based Annotation
In 3D modalities like CT and MRI, adjacent 2D slices often exhibit significant semantic redundancy. Consequently, annotating only the key slices of each sample can reduce annotation costs. AL works mentioned in this section select 2D slices within a 3D volume for annotation. Representativeness-based methods have been widely applied in this line of work. For instance, Zheng et al. (2020) utilized autoencoders to learn the semantic features of each slice, then selected and annotated key slices from axial, sagittal, and coronal planes with a strategy similar to RA (Zheng et al., 2019). Specifically, they initially trained three 2D segmentation networks and one 3D segmentation network, where the inputs for the 2D networks are slices from different planes. These segmentation networks were used to generate four sets of pseudo-labels and subsequently to train the final 3D segmentation network. Results showed that this slice-based strategy outperforms uniform sampling. Building upon this method, Peng et al. (2022) adopted a similar strategy in 3D knee cartilage and bone segmentation. Besides, Wu et al. (2022d) incorporated a self-attention module into the autoencoder to enhance slice-level feature learning.
Uncertainty methods have also been introduced for selecting key slices. Zhou et al. (2021b) introduced a quality assessment module to select slices with the highest predicted average IoU score. In muscle segmentation of CT images, Hiasa et al. (2020) selected key slices and key regions. This work adopted clustering to select key slices and further selected regions with high uncertainty within each key slice for annotation.
In recent years, hybrid strategies combining both uncertainty and representativeness were proposed for slice-based annotation. In shoulder MRI musculoskeletal segmentation, Ozdemir et al. (2021) adopted the variance of multiple MC dropout runs as the uncertainty metric. The posterior probability estimated by infoVAE (Zhao et al., 2017) was used as the representativeness metric. Li et al. (2024) proposed a hybrid strategy to select informative slices in musculoskeletal segmentation of lower extremities, where uncertainty was estimated with a Bayesian U-net while the representativeness was based on cosine similarity. They further adopted mutual information to minimize the sample redundancy following Nath et al. (2021). The proposed method achieved impressive performances on both the MRI and CT datasets.
@@ 5.2.2 One-shot Annotation
Currently, most AL works require multiple rounds of annotation. However, this setting could be impractical in medical image segmentation. Multi-round annotation requires physicians to be readily available for each round of labeling, which is unrealistic in practice. If physicians cannot complete the annotations on time, the AL process must be suspended. In contrast, one-shot annotation eliminates the need for multiple interactions with physicians. It also allows for selecting valuable samples in a single round, thus reducing time costs. Both one-shot annotation and cold-start AL aim to select the most optimal initial annotations. However, the former allows for a higher annotation budget and strictly limits the number of interactions with experts to just one. Most relevant works combine self-supervised features and specific sampling strategies to achieve one-shot annotation. For example, RA (Zheng et al., 2019) is one of the earliest works in one-shot AL for medical image segmentation. They applied the VAE feature and a representativeness strategy to select informative samples for annotation in one shot. RA performed excellently in gland segmentation of pathological images, whole-heart MRI images, and fungal of electron microscopic images. Wu et al. (2022d) proposed a representativeness-based framework for selecting key slices for annotation in one shot. They adopted self-learning to learn the semantic representation of each slice and used it to propagate the expert annotation to different slices. Jin et al. (2022b) combined features of contrastive learning with farthest-first sampling to achieve one-shot annotation. The proposed method demonstrated effectiveness on the ISIC 2018 and lung segmentation datasets. Additionally, Jin et al. (2023b) utilized auto-encoding transformations for self-supervised feature learning. They selected and annotated samples with high density based on reachable distance.
@@ 5.2.3 Annotation Cost
Current AL works often assume equal annotation costs for each sample. Yet, this is not the case in medical image segmentation, where the time to annotate different samples can differ greatly. AL techniques can better support physicians by considering annotation costs (e.g., annotation time). In detecting intracranial hemorrhage of CT scans, Kuo et al. (2018) combined predictive disagreement with annotation time to select samples for annotation. Specifically, they adopted the Jensen-Shannon divergence to measure the disagreement between the outputs of multiple models. Annotation time for each sample was estimated by the length of the segmentation boundary and the number of connected components. In this work, AL was framed as a 0-1 knapsack problem, and dynamic programming is used to solve this problem for selecting informative samples. In brain tumor segmentation, Shen et al. (2021) derived the annotation cost of a slice based on the distance between the queried slices and the already-labeled slices. Specifically, lower distance represented lower annotation cost. The rationale is that the annotation cost of labeling a similar slice would be cheaper than that of the unfamiliar slices. In brain structure segmentation, Atzeni et al. (2022) further considered the spatial relationships between multiple regions of interest to more accurately estimate the annotation cost. Moreover, the average Dice coefficient of previous rounds was used to predict the average Dice for current segmentation results. They selected and annotated regions that can maximize the average Dice.
5.2.4 Interactive Segmentation
Despite the success of automatic segmentation in medical imaging, there is still a potential for errors in clinical applications due to domain shifts or unseen ROIs. Interactive segmentation (Budd et al., 2021; Luo et al., 2021) could produce real-time adjustment of the current segmentations based on the user inputs of clicks, bounding boxes, or scribbles. As a result, interactive segmentation could rapidly tune the model towards current clinical applications with the guidance of doctors. For the sake of flexibility, current interactive segmentation methods accept annotations for any position. However, such a paradigm would be more efficient when the model itself could suggest where to annotate, which is exactly what active learning is good at. Therefore, combining AL and interactive segmentation would further reduce the annotation cost. In this section, all the mentioned papers worked interactively with different labeling units. Before the DL era, Su et al. (2015) had already integrated AL in the interactive cell segmentation. They selected the most informative superpixels for interactive annotation with expected prediction error. In MRI fetal brain segmentation, Wang et al. (2020a) proposed an uncertainty-guided framework for interactive refinement. They developed a novel network architecture to produce multiple segmentation results simultaneously, and the variance between different predictions served as the uncertainty metric. Slices with the highest uncertainty were fetched for interactive refinement by human experts. In interactive segmentation of 3D medical images, (Zhou et al., 2022) proposed a quality predictor, which produced a predicted IoU score with the current segmentation for each slice. With the interactive segmentation network, the quality predictor suggested slices with lower scores for expert annotation which could be in the form of scribble, bounding box, or extreme clicking. In Li et al. (2023), the most informative foreground and background superpixels were selected for interactive annotation.
5.3 Active Learning for Medical Image Reconstruction
AL can also be applied in medical image reconstruction. AL methods can help minimize the observations needed for modalities that require a long imaging time. This accelerates the imaging process and shortens the waiting period for patients. In this section, we’ll explore the application of AL in the reconstruction of MRI, CT, and electron microscopy. Please refer to Table. 4 for more detail.
@@ 5.3 Active Learning for Medical Image Reconstruction
In addition to MRI imaging, AL has been employed in CT reconstruction as illustrated by Wang et al. (2022a). They adaptively chose the scanning angles tailored to individual patients, leading to a reduction in both radiation exposure and scanning duration. In electron microscopy, Mi et al. (2020) initially enhanced low-resolution images to high-resolution and then predicted the location of region-of-interest and reconstruction error. A weighted DPP based on reconstruction error was applied to select pixels that needed to be rescanned. Results showed that weighted DPP maintained both low reconstruction error and spatial diversity.
Table 4: Summarization of surveyed works of active learning in medical image reconstruction.
Year Venues Modality ROIs Dataset Clinical Task
Jin et al. (2019) 2019 arXiv MRI Heart Cardiac Atlas Project MRI Reconstruction
Knee fastMRI MRI Reconstruction
Zhang et al. (2019) 2019 CVPR MRI Knee fastMRI MRI Reconstruction
Mi et al. (2020) 2020 MICCAI Electron Microscopy Mouse Cortex SNEMI3D Accelerated Acquisition of Electron Microscopy
Human Cerebrum in-house
Pineda et al. (2020) 2020 MICCAI MRI Knee fastMRI MRI Reconstruction
Bakker et al. (2020) 2020 NeurIPS MRI Knee fastMRI MRI Reconstruction
Brain fastMRI MRI Reconstruction
Wang et al. (2022a) 2022 arXiv CT Lung AAPM CT Reconstruction
Spine VerSe CT Reconstruction
6 Performance Evaluation of Active Learning in Medical Image Analysis
In the field of medical image analysis, there is now an increasing amount of AL works. Despite its rapid development, AL for medical image analysis still faces several issues that limit its application in real-world clinical tasks. On one hand, there is a lack of comprehensive evaluation of AL methods on the medical imaging datasets. Most AL works conducted experiments on standard datasets, such as CIFAR-10, CIFAR-100, or MedMNIST. However, real-world medical imaging datasets often contain less available data and higher complexity for analysis (Varoquaux and Cheplygina, 2022). Some AL works focused on a specific domain of medical imaging and achieved excellent performance, but their potential to generalize to a wider aspect of applications remains questionable. Beyond that, different AL methods exhibit inconsistent performance and may not necessarily outperform random sampling. In AL of classification tasks, Munjal et al. (2022) highlighted the absence of a consistently outperforming AL method and the fact that random sampling performs relatively well. In §4.2, we have also mentioned that the AL methods are inferior to random sampling when the annotation budget is low. As a result, we are uncertain about which AL method works according to our requirements and whether this method could outperform the most straightforward baseline, random sampling.
To clarify the aforementioned issues, we have conducted a comprehensive evaluation of different AL methods on multiple medical imaging datasets. The adopted three datasets are widely used by the entire medical imaging community. They also correspond to different modalities, organs, and tasks (e.g., classification and segmentation). We choose the most representative and popular AL methods for their evaluations of the medical imaging datasets. Besides, we provided the details of dataset splits, network architecture, and training hyperparameters for better reproducibility. Codes are also available on our accompanying website  [^4] https://github.com/LightersWang/Awesome-Active-Learning-for-Medical-Image-Analysis/tree/main/code.
6.1 Experimental Setups
Table 5: Training, validation, and testing splits of each dataset. Unless specified otherwise, the figures presented in this table represent the number of images of each split.
NCT-CRC-HE-100K ISIC 2020 ACDC
Training 90,000 20,869
656
Validation 10,000 2,319
10
Testing
7,180
(CRC-VAL-HE-7K)
9,938
20
6.1.1 Datasets
In this survey, we chose three medical imaging datasets for the performance evaluation of the AL methods, including two classification datasets and one segmentation dataset. The descriptions and dataset split are presented as follows, where a summarized table of the dataset split is in Table 5.
NCT-CRC-HE-100K (Kather et al., 2019): This dataset contains 100,000 patches from 86 hematoxylin & eosin (H&E) stained histological slides of human colorectal cancer and normal tissue. All the patches are $224\times 224$ at 0.5 microns per pixel. The patches are grouped into nine classes of different tissues, including adipose (ADI), background (BACK), debris (DEB), lymphocytes (LYM), mucus (MUC), smooth muscle (MUS), normal colon mucosa (NORM), cancer-associated stroma (STR), and colorectal adenocarcinoma epithelium (TUM). For the dataset split, we divided the dataset into training and validation sets with a ratio of 9:1 and utilized an additional dataset CRC-VAL-HE-7K which is provided by the same authors as the testing set. CRC-VAL-HE-7K shares the same acquisition protocol and tissue classes with NCT-CRC-HE-100K but contains 7,180 patches from 50 patients other than the patients of NCT-CRC-HE-100K.
ISIC 2020 (Rotemberg et al., 2021): ISIC 2020 is composed of 33,126 dermoscopic images from over 2,000 patients. Each image is labeled as benign or malignant by either doctors, long-term follow-up, or histopathology. ISIC 2020 contains 32,542 images of benign lesions and only 584 images of malignant lesions. We split this dataset of training, validation, and testing sets with a ratio of 6:1:3.
ACDC (Bernard et al., 2018): This dataset contains short-axis cardiac cine-MR images from 100 patients. In this survey, we only adopted the end-diastolic frame of each patient for evaluating different AL methods, which resulted in a total of 100 scans. Each scan corresponds to a human-annotated segmentation mask of the left ventricle (LV), myocardium (MYO), and right ventricle (RV). We followed the split from Luo et al. (2022), which contains 70, 10, and 20 scans in the training, validation, and testing sets, respectively. Due to the large spacing along the z-axis, 2D segmentation is more appropriate compared to 3D segmentation. Therefore, we trained the segmentation model with 2D slices and evaluated it with 3D volumes following Bai et al. (2017). Therefore, the training set is composed of 656 slices.
6.1.2 Evaluation Metrics
We employed different evaluation metrics for the task of each dataset. For the multi-class classification of NCT-CRC-HE-100K, we adopted the accuracy (ACC) to evaluate the classification performance. Due to the heavy class-imbalance of the binary classification task of ISIC 2020, we adopted the area under the receiver operating characteristic curve (AUC) for evaluation. For the segmentation task of ACDC, two well-known metrics of Dice similarity coefficient (DSC) and average surface distance (ASD) are used. DCS ranges from 0% (non-overlap) to 100% (perfect segmentation), and the lower ASD indicates a better alignment between the segmentation prediction and ground truth. To evaluate the overall performance, we presented the mean DSC and ASD of LV, MYO, and RV. Evaluation metrics on the testing set were reported as the final results.
Table 6: Accuracy of different active learning methods for multi-class pathological tissue classification. We reported the test performance of the initially labeled dataset and other active learning rounds with mean and standard deviation. The best and second-best results are bolded in red and blue, respectively.
ACC (%) 50 100 150 200 250 300
Random 60.46 64.39 $\pm$ 0.34 66.80 $\pm$ 3.66 71.77 $\pm$ 3.35 73.55 $\pm$ 2.38 75.87 $\pm$ 1.21
Confidence 63.61 $\pm$ 2.49 66.16 $\pm$ 3.12 66.29 $\pm$ 3.84 70.65 $\pm$ 2.85 71.69 $\pm$ 0.97
Entropy 62.65 $\pm$ 0.79 65.66 $\pm$ 1.31 68.78 $\pm$ 2.34 69.12 $\pm$ 2.99 68.86 $\pm$ 3.84
Margin 70.65 $\pm$ 2.20 71.33 $\pm$ 1.54 75.93 $\pm$ 3.00 74.64 $\pm$ 4.21 76.61 $\pm$ 2.90
BALD 62.13 $\pm$ 1.09 64.17 $\pm$ 4.20 69.34 $\pm$ 4.53 66.94 $\pm$ 2.50 68.49 $\pm$ 2.75
DBAL 61.73 $\pm$ 2.43 63.76 $\pm$ 5.33 67.8 $\pm$ 1.78 67.90 $\pm$ 3.06 71.48 $\pm$ 3.38
BADGE 68.52 $\pm$ 6.57 72.27 $\pm$ 3.45 74.69 $\pm$ 0.85 75.96 $\pm$ 2.65 75.74 $\pm$ 2.22
Core-Set-L2 62.73 $\pm$ 4.33 66.26 $\pm$ 2.44 64.72 $\pm$ 3.94 66.25 $\pm$ 2.98 64.66 $\pm$ 1.24
Core-Set-Cosine 63.07 $\pm$ 1.84 63.12 $\pm$ 4.20 69.32 $\pm$ 1.37 68.25 $\pm$ 2.30 68.21 $\pm$ 3.48
Fully supervised 93.36
(a) Low budget ($b=50$)
ACC (%) 1000 2000 3000 4000 5000 6000
Random 83.38 87.16 $\pm$ 1.43 87.68 $\pm$ 1.15 89.00 $\pm$ 0.91 89.04 $\pm$ 2.41 89.09 $\pm$ 1.08
Confidence 87.93 $\pm$ 0.76 87.46 $\pm$ 0.96 88.45 $\pm$ 0.64 90.01 $\pm$ 1.22 87.93 $\pm$ 3.04
Entropy 89.39 $\pm$ 0.23 86.54 $\pm$ 2.33 87.70 $\pm$ 2.25 88.04 $\pm$ 2.89 89.06 $\pm$ 1.57
Margin 88.60 $\pm$ 0.62 88.05 $\pm$ 1.83 88.13 $\pm$ 1.74 89.90 $\pm$ 0.63 86.89 $\pm$ 2.99
BALD 87.01 $\pm$ 0.99 85.85 $\pm$ 1.96 90.11 $\pm$ 0.53 89.19 $\pm$ 2.46 88.24 $\pm$ 2.21
DBAL 89.55 $\pm$ 0.75 86.56 $\pm$ 1.86 86.92 $\pm$ 3.62 88.76 $\pm$ 1.97 88.44 $\pm$ 0.53
BADGE 87.31 $\pm$ 0.51 87.56 $\pm$ 0.90 88.66 $\pm$ 2.33 88.39 $\pm$ 0.97 87.17 $\pm$ 1.30
Core-Set-L2 86.82 $\pm$ 0.37 88.99 $\pm$ 1.74 89.92 $\pm$ 0.98 89.34 $\pm$ 1.87 88.64 $\pm$ 3.10
Core-Set-Cosine 87.40 $\pm$ 0.49 88.09 $\pm$ 2.35 89.51 $\pm$ 1.20 89.80 $\pm$ 3.01 87.19 $\pm$ 2.78
Fully supervised 93.36
(b) High budget ($b=1000$)
6.1.3 Active Learning Settings
In this study, we performed $T=5$ rounds of annotation. To investigate how the number of queried samples in each round (i.e., annotation budget) affects the performance of different AL methods, we set different levels of annotation budgets $b$ for each dataset. Following Lüth et al. (2024), high ($b=1000$) and low ($b=50$) budgets were used for classification tasks of NCT-CRC-HE-100K and ISIC 2020. Following Gaillochet et al. (2023b), we adopted a budget of 10 slices ($b=10$) for ACDC segmentation. Considering the sizes of the involved datasets, the low budgets for classification ($b=50$) and segmentation ($b=10$) provide a chance to look into the performance of different AL methods in the low data regime. Before the active learning process started, we randomly selected the initial labeled pool to train an initial model. The size of the initial pool is equal to the annotation budget. The model training and sample selection are seeded. We ran each active learning method of a specific budget and dataset five times with different random seeds and reported the mean and standard deviation as the result.
6.1.4 Comparison Methods
For fairness and reproducibility, we conducted the evaluation using the following methods: Random: the baseline of active learning, which randomly draws the unlabeled samples. Confidence, Entropy, and Margin Lewis and Catlett (1994); Joshi et al. (2009); Roth and Small (2006): these methods are all classic uncertainty-based AL methods, which calculated the confidence, entropy, and margin with the prediction probability as the uncertainty scores. Lower confidence, higher entropy, and lower margin indicate higher uncertainty. DBAL Gal et al. (2017): This method integrated entropy and MC dropout for better uncertainty estimation. During sample selection, the model runs multiple times with all dropout layers activated, and the average probability of all MC dropout runs is used for calculating entropy. BALD Gal et al. (2017): This method calculated BALD as the uncertainty score, which aims to maximize the mutual information between predictions and model parameters. MC dropout was also used in this method. Core-Set Sener and Savarese (2018): This method performed cover-based sampling using the feature embedding of each sample. For the balance of computation time and performance, we used the k-Center-Greedy for sample selection. To investigate how the distance metric affected the AL performance, we proposed a variant of Core-Set named “Core-Set-Cosine” which replaced the original L2 distance with the cosine distance. The original Core-Set was referred to as “Core-Set-L2” to avoid confusion. BADGE Ash et al. (2020): This method applied gradient as uncertainty estimation and utilized k-Means++ to improve diversity. Specifically, the gradient of cross-entropy loss was used in the classification task while the segmentation task used the gradient of the sum of the Dice loss and cross-entropy loss.
It should be noted that uncertainty-based methods in segmentation are slightly different from those of the classification. Specifically, we first produced the pixel-wise scores and then utilized the averaged scores for sample selection in the segmentation task.
Table 7: AUC of different active learning methods for the binary skin lesion classification. We reported the test performance on ISIC 2020 of the initially labeled dataset and other active learning rounds with mean and standard deviation. The best and second-best results are bolded in red and blue, respectively.
AUC 50 100 150 200 250 300
Random 0.5609 ± 0.0325 0.5408 ± 0.0228 0.5807 ± 0.0391 0.5679 ± 0.0632 0.6143 ± 0.0610
Confidence 0.5432 ± 0.0233 0.5713 ± 0.0472 0.5622 ± 0.0302 0.5584 ± 0.0171 0.5649 ± 0.0794
Entropy 0.5375 ± 0.0182 0.5432 ± 0.0209 0.5919 ± 0.0413 0.5537 ± 0.0383 0.5543 ± 0.0220
BALD 0.5865 ± 0.0235 0.5828 ± 0.0499 0.5818 ± 0.0800 0.5977 ± 0.0273 0.5682 ± 0.0313
DBAL 0.5322 ± 0.0300 0.5344 ± 0.0226 0.5601 ± 0.0341 0.5879 ± 0.0274 0.5442 ± 0.0350
BADGE 0.5601 ± 0.0373 0.5648 ± 0.0333 0.6219 ± 0.0757 0.6246 ± 0.0894 0.6172 ± 0.0890
Core-Set-L2 0.5834 ± 0.0344 0.6761 ± 0.0382 0.5956 ± 0.0610 0.5934 ± 0.0464 0.6313 ± 0.0867
Core-Set-Cosine 0.574 0.5459 ± 0.0367 0.6361 ± 0.0621 0.6032 ± 0.0402 0.6127 ± 0.0586 0.6317 ± 0.0797
Fully supervised 0.8424
(a) Low budget ($b=50$)
AUC 1000 2000 3000 4000 5000 6000
Random 0.6898 ± 0.0216 0.7258 ± 0.0249 0.7556 ± 0.0236 0.7551 ± 0.0157 0.7718 ± 0.0124
Confidence 0.7058 ± 0.0120 0.7525 ± 0.0069 0.7760 ± 0.0171 0.7957 ± 0.0184 0.7949 ± 0.0113
Entropy 0.7063 ± 0.0144 0.7478 ± 0.0090 0.7692 ± 0.0117 0.7838 ± 0.0062 0.7875 ± 0.0101
BALD 0.6851 ± 0.0156 0.7489 ± 0.0144 0.7683 ± 0.0097 0.7947 ± 0.0098 0.7837 ± 0.0073
DBAL 0.7120 ± 0.0195 0.7613 ± 0.0053 0.7616 ± 0.0261 0.7896 ± 0.0120 0.7883 ± 0.0127
BADGE 0.7074 ± 0.0264 0.7456 ± 0.0103 0.7734 ± 0.0150 0.7983 ± 0.0100 0.7917 ± 0.0064
Core-Set-L2 0.7304 ± 0.0303 0.7511 ± 0.0156 0.7499 ± 0.0106 0.7588 ± 0.0056 0.7575 ± 0.0067
Core-Set-Cosine 0.644 0.7288 ± 0.0236 0.7593 ± 0.0177 0.7807 ± 0.0182 0.7925 ± 0.0135 0.7917 ± 0.0109
Fully supervised 0.8424
(b) High budget ($b=1000$)
6.1.5 Implementation Details
Classification: For all the classification task, we used ResNet-18 (He et al., 2016) as the backbone and the loss function is cross-entropy. We trained the model using stochastic gradient descent with momentum for 100 epochs with a batch size of 128. The learning rate and momentum were set as 0.01 and 0.9, respectively. Also, the cosine learning rate decay was adopted for smoother convergence. Data augmentations of the input image are different in the two classification datasets. In NCT-CRC-HE-100K, we only used the random horizontal flip. For ISIC 2020, we followed the data augmentation from Zhuang et al. (2018) which includes random crop, flip, rotation, affine transforms, and color jittering.
Segmentation: We used a 5-level U-Net (Ronneberger et al., 2015) for segmentation. Each level of the encoder or decoder contains two blocks. Each block consists of a 2D convolution, dropout layer with a probability of 0.1, batch normalization, and leaky ReLU activation. The segmentation loss is the combination of cross-entropy loss and Dice loss. We trained the model using the Adam optimizer (Kingma and Ba, 2014) for 4,000 iterations with a batch size of 32. The learning rate is 0.001 while decaying along the training iterations with the polynomial schedule. Data augmentation includes random flip, rotation of 90 degrees, and rotation of arbitrary degrees.
Experiments were conducted on NVIDIA GeForce RTX 3090 and 4090 GPUs and the CUDA version is 11.3. Codes are implemented using Python (version 3.8.10) and the PyTorch framework (version 1.11.0).
6.2 Experimental Results and Performance Analysis
6.2.1 Active Learning Results for Pathological Tissue Classification
We first evaluated the active learning performance on the pathological tissue classification task. The results of the testing accuracy are shown in Table 6. Margin performed well in the low-budget scenario. The reason for that is this method exploits the information of the wrong predictions of a similar class, which is in line with the finding in Hu et al. (2023). BADGE performs well in the low-budget scenarios largely due to the k-Means++ clustering. However, its performance drops in the high-budget scenario, which may indicate that the gradient embeddings are less suitable in AL when there is a distribution shift between the training and testing sets. The results of this section call for a more in-depth investigation of the generalizability to distribution shift of the AL methods.
6.2.2 Active Learning Results for Skin Lesion Classification
We have also conducted thorough evaluations on the ISIC 2020 dataset, which corresponds to a binary classification problem with severe class-imbalance. The AUC of the test split is shown in Table 7. It should be noted that Confidence and Margin are equivalent in the binary setting, so we only report the results of the former. In the low-budget scenario, Core-Set and its variant achieved better performances compared to the uncertainty-based methods. It indicates that representativeness-based methods or methods with improved diversity are more favored than uncertainty-based ones when the budget is low and the task is extremely difficult. Among all the uncertainty-based methods, BADGE stands out in certain rounds for its clustering operations that enhance diversity. For the high budget, the performance of the Core-Set variants is still competitive. However, the performance of the uncertainty-based methods improved. Results here demonstrated how the annotation budget affects the performance of the uncertainty-based and representativeness-based methods, in which the former fits a higher budget while the latter fits a lower budget.
Table 8: Mean DSC and ASD of different active learning methods for cardiac MRI segmentation. We reported the test metrics on ACDC of the initially labeled dataset and other active learning rounds with mean and standard deviation. The best and second-best results are bolded in red and blue, respectively.
DSC 10 20 30 40 50 60
Random 0.7872 ± 0.0552 0.8285 ± 0.0379 0.8657 ± 0.0063 0.8656 ± 0.0089 0.8721 ± 0.0047
Confidence 0.6805 ± 0.0205 0.7499 ± 0.0336 0.8347 ± 0.0203 0.8696 ± 0.0057 0.8717 ± 0.0065
Entropy 0.6888 ± 0.0088 0.7712 ± 0.0393 0.8538 ± 0.0117 0.8639 ± 0.0062 0.8714 ± 0.0040
Margin 0.6841 ± 0.0124 0.7360 ± 0.0338 0.8214 ± 0.0535 0.8640 ± 0.0096 0.8734 ± 0.0053
BADGE 0.8192 ± 0.0144 0.8486 ± 0.0151 0.8632 ± 0.0060 0.8731 ± 0.0032 0.8698 ± 0.0080
Core-Set-L2 0.4966 0.8112 ± 0.0210 0.8491 ± 0.0144 0.8570 ± 0.0074 0.8572 ± 0.0057 0.8660 ± 0.0073
Core-Set-Cosine 0.7344 ± 0.0234 0.8248 ± 0.0109 0.8458 ± 0.0225 0.8549 ± 0.0204 0.8678 ± 0.0089
Fully supervised 0.9045
(a) Mean DSC
ASD (mm) 10 20 30 40 50 60
Random 4.71 ± 1.13 2.82 ± 0.37 3.13 ± 0.48 2.05 ± 0.70 2.21 ± 0.73
Confidence 6.63 ± 2.05 3.60 ± 1.39 2.86 ± 0.39 2.21 ± 0.98 2.22 ± 0.62
Entropy 7.80 ± 1.70 3.27 ± 1.22 3.00 ± 1.08 2.24 ± 1.23 2.67 ± 0.84
Margin 6.64 ± 3.17 4.65 ± 1.94 2.66 ± 0.76 3.08 ± 1.02 3.22 ± 1.41
BADGE 4.22 ± 1.10 4.14 ± 0.32 3.10 ± 0.70 2.62 ± 0.81 2.56 ± 0.97
Core-Set-L2 11.36 3.22 ± 0.69 2.84 ± 0.66 2.71 ± 0.76 3.40 ± 1.31 2.86 ± 0.76
Core-Set-Cosine 4.75 ± 1.57 2.74 ± 0.52 3.50 ± 1.36 2.33 ± 0.88 1.85 ± 0.31
Fully supervised 2.09
(b) Mean ASD
6.2.3 Active Learning Results for MRI Cardiac Segmentation
For segmentation, we evaluated different AL methods on the ACDC dataset. We reported the mean DSC and ASD of the segmentation results in Table 8. BADGE achieved the best or second-best performance in mean DSC for multiple rounds. Core-Set performed well on both the mean DSC and ASD in the early rounds of the lower budget scenario. Both two methods improve sampling diversity to some extent. However, for the later rounds, the performance of the uncertainty-based methods and random sampling improves in mean DSC and ASD. This result aligns with the findings in the previous section.
6.2.4 Effectiveness on Different Distances between Images
Distance measurement plays an important role in AL which may significantly impact the performance of AL algorithms. In this section, we evaluated the performance of the two most popular distances in AL, which are L2 and cosine distances. These two distances are based on the feature embedding. Assume $x$ stands for the sample itself and its corresponding feature embedding is $\mathbf{z}=\left[z_{1},z_{2},\ldots,z_{d}\right]$, $d$ is the feature dimension. Based on the feature embedding, the L2 distance between two images $x^{a}$ and $x^{b}$ is as follow:
[…]
while the cosine distance is:
$\text{Cosine}(x^{a},x^{b})=1-\frac{\mathbf{z}^{a}\cdot\mathbf{z}^{b}}{\|\mathbf{z}^{a}\|\cdot\|\mathbf{z}^{b}\|}=1-\frac{\sum_{i=1}^{d}z^{a}_{i}z^{b}_{i}}{\sqrt{\sum_{i=1}^{d}{z^{a}_{i}}^{2}}\sqrt{\sum_{i=1}^{d}{z^{b}_{i}}^{2}}}$ (13)
To conduct experiments, we replaced the L2 distance in Core-Set with the cosine distance.
Performance comparisons between Core-Set-L2 and Core-Set-Cosine are illustrated in Fig.6. Results on the NCT-CRC-HE-100K dataset showed no significant difference between the L2 and cosine distances across all budget levels. In ISIC 2020, the L2 distance tends to be better in the early rounds, indicating its ability to rapidly start the model, while Core-Set-Cosine significantly outperformed the Core-Set-L2 when the budget was high. On the ACDC dataset, Core-Set-L2 also outperformed Core-Set-Cosine in the early rounds but their performance after selecting more samples is similar. These results indicate that distance metrics play an important role in the performance of AL methods used in medical image analysis, and they should be carefully chosen according to the target tasks and the budgets. Generally speaking, L2 distance is more suitable for the low-budget scenario while the cosine distance might be a better choice when the budget is high.
(a) NCT-CRC-HE-100K ($b=1000$)
(b) NCT-CRC-HE-100K ($b=50$)
(c) ISIC 2020 ($b=1000$)
(d) ISIC 2020 ($b=50$)
(e) ACDC (Mean DSC)
(f) ACDC (Mean ASD)
Figure 6: Performance comparisons between the L2 and cosine distances on all three datasets. We used Core-Set as the base AL method.